OpenStudy (anonymous):

when the radius is 8 meters,What is the arc length inscribed by a 1500 angle? Round to the nearest tenth of a meter.

OpenStudy (anonymous):

A) 150 π m B) 8 π m C) 15.9 m D) 20.9 m

hartnn (hartnn):

Arc length is just \(\Large A= r \times \theta\) where r is the radius, but \(\theta \) is the central angle in RADIANS. so, angle needs to be converted from degrees to radians first.

hartnn (hartnn):

do u know how to convert 150 degrees into radians ?

OpenStudy (anonymous):

yes , 150 degrees will be 2.617993878 in radias

OpenStudy (anonymous):

so A= 8 x 2.617993878 = 20.94395102 so when the answer is rounded to the nearest tenth it will be 20.9 am i correct

hartnn (hartnn):

yes! thats correct :)

OpenStudy (anonymous):

thank you so much for your help @hartnn :)

hartnn (hartnn):

you're most welcome jody :)